OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62? Answer A. x = 6 B. x = –10 C. x = –6 D. x = 10

OpenStudy (anonymous):

= -24 / 2(2) = -6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

y=2(x^2+12x+31) =2((x+6)^2-5) so this will symmetric at x=-6